### v2

# HyperGaussians: High-Dimensional Gaussian Splatting for High-Fidelity Animatable Face Avatars

Abstract
We introduce HyperGaussians, a novel extension of 3D Gaussian Splatting for high-quality animatable face avatars. Creating such detailed face avatars from videos is a challenging problem and has numerous applications in augmented and virtual reality. While tremendous successes have been achieved for static faces, animatable avatars from monocular videos still fall in the uncanny valley. The de facto standard, 3D Gaussian Splatting (3DGS), represents a face through a collection of 3D Gaussian primitives. 3DGS excels at rendering static faces, but the state-of-the-art still struggles with nonlinear deformations, complex lighting effects, and fine details. While most related works focus on predicting better Gaussian parameters from expression codes, we rethink the 3D Gaussian representation itself and how to make it more expressive. Our insights lead to a novel extension of 3D Gaussians to high-dimensional multivariate Gaussians, dubbed ’HyperGaussians’. The higher dimensionality increases expressivity through conditioning on a learnable local embedding. However, splatting HyperGaussians is computationally expensive because it requires inverting a high-dimensional covariance matrix. We solve this by reparameterizing the covariance matrix, dubbed the ’inverse covariance trick’. This trick boosts the efficiency so that HyperGaussians can be seamlessly integrated into existing models. To demonstrate this, we plug in HyperGaussians into the state-of-the-art in fast monocular face avatars: FlashAvatar. Our evaluation on 19 subjects from 5 face datasets shows that HyperGaussians outperform 3DGS numerically and visually, particularly for high-frequency details like eyeglass frames, teeth, complex facial movements, and specular reflections.
Figure 1: We propose a novel representation for monocular face avatars: HyperGaussians. HyperGaussians extend 3D Gaussians to higher dimensions, resulting in improved high-frequency details for specular reflections and thin structures. HyperGaussians can be plugged into existing models with minimal overhead. This figure shows the effect of plugging HyperGaussian (bottom) into the state-of-the-art for monocular face avatars, FlashAvatar [51] (top). Note the improvement in specular reflections, complex deformations, and thin structures.
Figure 2: We propose an extension to 3D Gaussians, dubbed HyperGaussians, and plug them into an existing method for face avatars, FlashAvatar [51]. FlashAvatar modulates 3D Gaussian primitives with expression-dependent offsets $\Delta$. We make a single modification to the pipeline: We plug in HyperGaussians (Sec. 3.2) between the MLP output and the rasterization, which modifies the offsets $\Delta$ in higher dimensions. Instead of directly predicting offsets $\Delta$, we predict a latent $z_{\psi}$ that conditions HyperGaussians. Without any other modifications or hyperparameter tuning, this simple change enhances details in the final avatar (Fig. 7) and leads to a performance boost (Tab. 2). This figure has been adapted from FlashAvatar [51].
1 Introduction
Modelling dynamic scenes like human faces is a long-standing problem in 3D Computer Vision and has a variety of applications in augmented and virtual reality, the movie and gaming industry, and virtual telepresence [20, 36]. Faces exhibit highly nonlinear deformations with topology changes, such as eye blinks and mouth opening. In addition, faces show complex lighting effects like specular reflections in the eyes and on glasses [23, 24, 43, 45]. It is important to solve these details and cross the uncanny valley, enhancing human-avatar interaction, and creating truly lifelike avatars [35, 17]. One crucial element towards this goal is the development of high-quality and efficient 3D representations.
Recently, 3D Gaussian Splatting [18] has become the de facto standard for modeling humans due to its quality and efficiency [46, 8, 51, 63, 43]. The state-of-the-art for human faces rigs Gaussians with a Morphable Face Models [2, 13, 27]. Each Gaussian is attached to a mesh triangle and follows its deformation, for example, using linear blend-skinning [27, 29, 22]. Such linear models are a good approximation, but they cannot represent nonlinear deformations and specular effects. To mitigate this, recent works leverage Neural Networks to predict offsets based on an input expression [51, 46, 8]. This improves results over static Gaussians, but it is not sufficient to model details. Fig. 7 shows how these methods struggle to represent thin structures like teeth, hair, and glass frames (rows 1 - 4) and specular reflections on eyes and glasses (rows 1 and 3), and nonlinear deformations like closing eyelids (rows 2 and 4). These high-frequency details are key for crossing the uncanny valley.
In this work, we extend 3D Gaussian Splatting [18] to arbitrary higher dimensions and call the novel representation HyperGaussians (Sec. 3.2). HyperGaussians are an expressive, lightweight, and fast high-dimensional representation for high-quality face avatars. At their core, they are multivariate Gaussians. The additional dimensions and the local embeddings improve complex local deformations and specular effects. However, due to the higher dimensionality, a naïve implementation of high-dimensional Gaussians requires extremely high computational resources for splatting (Fig. 5). To solve this, we propose a re-parameterization, called the inverse covariance trick (Sec. 3.2). With the inverse covariance trick, high-dimensional Gaussian primitives can be splatted efficiently in real-time.
To the best of our knowledge, we are the first to improve high-frequency details in facial avatars by enhancing the low-level parameterization of the Gaussian primitives. Most related works focus on improving a model architecture or training scheme [51, 8, 46, 42, 21] that estimates the dynamics of the Gaussian parameters using neural networks. Our contribution, HyperGaussians, is orthogonal to these efforts. HyperGaussian builds on top of the 3DGS framework and thus can be readily plugged into an existing pipeline and boost its performance out-of-the-box, without any other changes to the model, architecture, or hyperparameters. We demonstrate this by integrating our proposed HyperGaussians into the state-of-the-art in fast face avatar learning from monocular videos, FlashAvatar [51]. The only change we make is to replace the 3D Gaussians with HyperGaussians, the rest of the model remains exactly the same. We evaluate on videos of $19$ subjects from $4$ different datasets. The result is a boost in performance for self- and cross-reenactment, as demonstrated in Figs. 1, 7, 8 and 3 and Tab. 2. HyperGaussian can model thin structures like glass frames and teeth, specular reflections on eyes and glasses, and complex nonlinear deformations such as closing of the eyelid.
In summary, this paper contributes the following:
  • A novel expressive representation for modeling high-frequency and dynamic effects, dubbed HyperGaussians.
  • The inverse covariance trick, a technical contribution that dramatically increases the computational efficiency of high-dimensional Gaussian conditioning.
  • A case study on $19$ subjects from $4$ datatasets, where we plug in HyperGaussian into the state-of-the-art fast face avatar, FlashAvatar, demonstrating improved rendering quality without any other changes to the model.
FlashAvatar
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
5 seconds 20 seconds 60 seconds 180 seconds 600 seconds
Figure 3: We plug in our proposed HyperGaussian into FlashAvatar [51] and compare the convergence speed. The only difference between FlashAvatar and Ours is the substitution of 3D Gaussians (top) with HyperGaussians (bottom), as described in Sec. 3.3. HyperGaussians displays sharper results throughout the training.
2 Related Work
Dynamic 3D Scenes
Traditional 3D representations, such as meshes and point clouds, are increasingly being replaced by more innovative methods. Neural Radiance Fields (NeRFs) [33], Neural Implicit Functions [37, 32], and 3D Gaussians [18] can represent 3D scenes with a high level of detail and enable photorealistic novel view synthesis. Neural Radiance Fields and Implicit Functions can be adapted to capture dynamic effects by conditioning on a time dimension [38, 39, 41] to enable the replay of videos from novel views and featuring moving objects or a change in facial expressions. NeRFies [38] optimizes a continuous deformation field that maps from a deformed to a canonical space but suffers from artifacts when topological changes arise, e.g., a mouth opening. HyperNeRF [39] showed that these issues can be mitigated by modeling the deformation field in higher dimensions. The HyperNeRF deformation field slices the high-dimensional deformation field with learnable embeddings. Inspired by HyperNeRF, we propose an extension to 3D Gaussian Splatting by modeling Gaussians in a higher-dimensional space.
N-dimensional Gaussian Splatting
A number of works have explored Gaussian primitives with non-3D dimensionalities [16, 50, 10, 15, 60, 30, 56]. 1D Cylindrical Gaussian Splatting is being used for hair modeling [60, 30, 56]. 2D Gaussian Splatting [16] uses oriented planar Gaussian disks to improve geometric consistency in Gaussian Splatting. 4D Gaussian Splatting [50] extends 3DGS for rendering dynamic scenes. They propose a custom CUDA kernel for splatting 4D Gaussians with a time dimension. This approach reconstructs detailed videos, but it does not support higher dimensions. NDGS [10] formulates a Gaussian Mixture Model for representing static scenes with high appearance variability, which is effective for highly reflective surfaces. Their representation exhibits some key differences from ours. Their representation has no degrees of freedom for the conditional orientation of the 3D Gaussians, i.e., the Gaussians cannot rotate. Moreover, the rendered size of a Gaussian depends on the probability density function of the joint distribution. This causes Gaussians to disappear under large deformations and leads to semantic subparts of the scene being modeled by multiple Gaussians, which are invisible most of the time. Our formulation removes these limitations and ensures that the Gaussians deform consistently.
Face Avatars
A popular goal is reconstructing a face avatar with a high level of fidelity and rendering it under a novel pose and expression [25, 6, 31, 9, 26, 3, 5, 4, 58]. 3D Gaussians have become the de facto standard for modeling such face avatars [52, 25, 44, 31, 49, 54, 48, 59]. Tab. 1 provides an overview of the most closely related works. The state-of-the-art [51, 46, 21, 42] attaches Gaussians to the mesh of a Morphable Model (3DMM) [13, 2, 27]. 3DMMs serve as a strong shape prior and were already extensively made use of in implicit avatars, such as NerFACE [11] and INSTA [61]. This enables driving the avatar with controlled expressions and head poses with linear blend skinning [22, 27]. SplattingAvatar [46] improves over vanilla blendskinning by optimizing embeddings on the mesh. However, SplattingAvatar does not consider that deformations also affect the appearance, for example, a head pose change will move the specular reflections on glasses (Fig. 7), leading to a blurry rendering. Recent works take a step towards improving such high-frequency effects by predicting expression-dependent Gaussian parameter offsets [51, 63, 9, 8], allowing the Gaussian parameters to change for different expressions. Predicting these expression-dependent offsets makes FlashAvatar [51] the current state-of-the-art for fast face avatar modeling.
Method Representation Local context Dynamic context
GaussianAvatars [42] 3DGS $\times$ $\times$
SurFHead [21] 2DGS $\times$ $\times$
SplattingAvatar [46] 3DGS $\checkmark$ $\times$
MonoGaussianAvatar [8] 3DGS $\times$ $\checkmark$
FlashAvatar [51] 3DGS $\times$ $\checkmark$
FlashAvatar w/ HyperGaussian (Ours) HyperGaussian $\checkmark$ $\checkmark$
Table 1: We show the differences to the most closely related works. GaussianAvatars [42] and SuRFHead [21] deform 3D Gaussians based on an underlying FLAME mesh [27] without local embeddings or dynamic inputs like facial expressions. SplattingAvatar [46] optimizes local embeddings but the Gaussian properties are not dependent on expressions or pose. MonoGaussianAvatar [8] and FlashAvatar [51] predict expression-dependent offsets to the Gaussian properties, but their lack of local context leads to blurry or distorted results, see Fig. 7. Our proposed representation attaches high-dimensional Gaussians to the mesh and optimizes learnable local embeddings for modulating the Gaussian properties based on expressions.
3 Method
We propose a novel representation for modeling dynamic 3D scenes and apply it to face avatars. Our novel representation extends 3D Gaussians [18] to higher dimensions and extends each Gaussian primitive with an embedding that provides local context. This enables representing finer details such as a thin frame of glasses, tiny gaps between teeth, specular reflections on the eyes, and glasses (Fig. 1). We dub our novel representation HyperGaussians.
We start with preliminaries in Sec. 3.1 and introduce the general framework of HyperGaussians in Sec. 3.2. While already performing well for a low-dimensional hyperspace (less than 5 hyper-space dimensions, Tab. 3), a naïve implementation does not scale because of higher computational and memory load (Fig. 5). As a solution, we propose the inverse covariance trick. We demonstrate the effectiveness of HyperGaussians to reconstruct detailed face avatars from monocular videos in Sec. 3.3. Specifically, we plug in HyperGaussians into an existing method, FlashAvatar [51]. A simple replacement of 3D Gaussians with HyperGaussians without any hyperparameter-tuning boosts FlashAvatar’s performance (Tab. 2).
3.1 Preliminary: 3D Gaussian Splatting
3D Gaussian Splatting [18] (3DGS) models a static scene with colored anisotropic 3D Gaussians. The Gaussians are parameterized by their mean ${\boldsymbol{\mu}\in\mathbb{R}^{3}}$ and covariance matrix ${\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}}$ and their appearance is represented by their opacity $\alpha$ and a color $\boldsymbol{c}$, where common choices are RGB color or Spherical Harmonics for view-dependent effects. Given images, camera parameters, and a sparse estimated point cloud, 3DGS can reconstruct a scene by optimizing a set of Gaussians via differentiable rasterization.
To ensure that the covariance matrices remain positive semi-definite during optimization, 3DGS [18] first defines a parametric ellipsoid using a scaling matrix $\boldsymbol{S}$ and a rotation matrix $\boldsymbol{R}$, then constructs the covariance matrix as
$\boldsymbol{\Sigma}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{\top}\boldsymbol{R}^{\top}.$ (1)
These matrices are themselves parameterized by a scaling vector ${\boldsymbol{s}\in\mathbb{R}^{3}}$ and a unit quaternion ${\boldsymbol{q}\in\mathbb{R}^{4}}$, respectively. We denote the set of Gaussians as $\{(\boldsymbol{\mu}_{i},\boldsymbol{s}_{i},\boldsymbol{q}_{i},\boldsymbol{c}_{i},\alpha_{i})\}_{i=1}^{n}$.
Novel views can be rendered by splatting the Gaussians, following these two steps: the Gaussians are first projected onto the camera plane and then alpha blended together with respect to their evaluated density $G(\boldsymbol{x})=\exp\bigl{(}-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})\bigr{)}$ and learned opacity $\alpha$. For each pixel, its color can then be computed using alpha blending:
$\boldsymbol{C}=\sum_{i}\boldsymbol{c}_{i}^{\prime}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}).$ (2)
3.2 HyperGaussians
This section introduces HyperGaussians, our proposed replacement for vanilla 3D Gaussians. HyperGaussians are an extension of 3D Gaussians to higher dimensions. As an example, a vanilla Gaussian in Kerbl et al. [18] has $m=3$ dimensions for the mean $\boldsymbol{\mu}$ representing its position. Intuitively, our HyperGaussian generalizes the vanilla Gaussian primitive to $(m+n)$ dimensions. We call $m$ the attribute dimensionality and the additional dimensions $n$ the latent dimensionality.
Formulation
This paragraph describes the HyperGaussian more formally. Consider a random vector $\boldsymbol{\gamma}\sim\mathcal{N}\bigl{(}\boldsymbol{\mu},\boldsymbol{\Sigma}\bigr{)}$ that is partitioned into vectors $\boldsymbol{\gamma}=(\boldsymbol{\gamma}_{a},\boldsymbol{\gamma}_{b})^{\top}$, where $\boldsymbol{\gamma}_{a}\in\mathbb{R}^{m}$ and $\boldsymbol{\gamma}_{b}\in\mathbb{R}^{n}$.
The partitioning of $\boldsymbol{\gamma}$ leads to the following block matrix view of $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$:
$\boldsymbol{\mu}=\begin{bmatrix}\boldsymbol{\mu}_{a}\\ \boldsymbol{\mu}_{b}\end{bmatrix},\quad\boldsymbol{\Sigma}=\begin{bmatrix}\boldsymbol{\Sigma}_{aa}&\boldsymbol{\Sigma}_{ab}\\ \boldsymbol{\Sigma}_{ba}&\boldsymbol{\Sigma}_{bb}\end{bmatrix},$ (3)
with $\boldsymbol{\Sigma}_{ba}=\boldsymbol{\Sigma}_{ab}^{\top}$. In practice, $\boldsymbol{\gamma}_{a}$ is instantiated with 3D Gaussian attributes (e.g., position) and $\boldsymbol{\gamma}_{b}$ with a latent code. Following NDGS [10], we parameterize the mean $\boldsymbol{\mu}$ directly and decompose the covariance $\boldsymbol{\Sigma}$ into its Cholesky factor $\boldsymbol{L}$, such that $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}$, where $\boldsymbol{L}$ is a lower triangular matrix with positive diagonal entries. Training requires optimizing the parameters of the Cholesky factor $\boldsymbol{L}$ together with the mean $\boldsymbol{\mu}$. Please read on to the end of this section for a detailed explanation of which parameters require training.
Splatting
HyperGaussians can be splatted with the differentiable rasterizer proposed by Kerbl et al. [18] after being reduced to the attribute dimensionality. This can be done by conditioning on the latent dimensions. Geometrically, conditioning corresponds to taking an $m$-dimensional slice through the multivariate Gaussian with $(m+n)$ total dimensions. Fig. 4 shows two examples for $m=n=1$. More formally, we are interested in the conditional distribution $p(\boldsymbol{\gamma}_{a}|\boldsymbol{\gamma}_{b})=\mathcal{N}\bigl{(}\boldsymbol{\mu}_{a|b},\boldsymbol{\Sigma}_{a|b}\bigr{)}$. Since our HyperGaussian follows a multivariate Gaussian distribution, the conditional distribution can be computed in closed form [1]:
$\begin{split}\boldsymbol{\mu}_{a|b}&=\boldsymbol{\mu}_{a}+\boldsymbol{\Sigma}_{ab}\boldsymbol{\Sigma}_{bb}^{-1}(\boldsymbol{\gamma}_{b}-\boldsymbol{\mu}_{b})\\ \boldsymbol{\Sigma}_{a|b}&=\boldsymbol{\Sigma}_{aa}-\boldsymbol{\Sigma}_{ab}\boldsymbol{\Sigma}_{bb}^{-1}\boldsymbol{\Sigma}_{ba}.\end{split}$ (4)
The output of this computation is then used to recover a vanilla 3D Gaussian and can be splatted with the differentiable rasterizer introduced by Kerbl et al. [18].
Figure 4: Gaussian Conditioning on two examples with large (left) and small (right) uncertainty at different realizations of $\boldsymbol{\gamma}_{b}$. As a result, the conditional mean shifts, while the conditional covariance is the same for both slices.
Inverse Covariance Trick for Fast Conditioning
A naïve implementation of the conditioning in Eq. 4 is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$. The bottleneck lies in storing and inverting the conditional covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$. The latent dimensionality ($n$) is typically much larger than the dimensionality after conditioning $m$ ($n\gg m$) and therefore expensive to invert.
To save the cost of this inversion, we apply the inverse covariance trick. The key idea is to reformulate the HyperGaussians in terms of their precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ such that $\boldsymbol{\gamma}\sim\mathcal{N}\bigl{(}\boldsymbol{\mu},\boldsymbol{\Lambda}^{-1}\bigr{)}$. Let’s consider the general block matrix view (as in Eq. 3)
$\boldsymbol{\Sigma}^{-1}=\boldsymbol{\Lambda}=\begin{bmatrix}\boldsymbol{\Lambda}_{aa}&\boldsymbol{\Lambda}_{ab}\\ \boldsymbol{\Lambda}_{ba}&\boldsymbol{\Lambda}_{bb}\end{bmatrix}$ (5)
with $\boldsymbol{\Lambda}_{ba}=\boldsymbol{\Lambda}_{ab}^{\top}$. The conditional mean and covariance matrix can now be expressed as
$\begin{split}\boldsymbol{\mu}_{a|b}&=\boldsymbol{\mu}_{a}-\boldsymbol{\Lambda}_{aa}^{-1}\boldsymbol{\Lambda}_{ab}\bigl{(}\boldsymbol{\gamma}_{b}-\boldsymbol{\mu}_{b}\bigr{)}\\ \boldsymbol{\Sigma}_{a|b}&=\boldsymbol{\Lambda}_{aa}^{-1}.\end{split}$ (6)
This new formulation only requires storing and inverting the much smaller block $\boldsymbol{\Lambda}_{aa}$ of the precision matrix. While this change seems minor in the derivation, it heavily improves both the speed and memory during runtime.
Optimizable Parameters
Eq. 6 is also helpful for understanding which parameters are optimized during training. It is not necessary to optimize the full precision matrix $\boldsymbol{\Lambda}$. The HyperGaussian parameters that require optimization are the means $\boldsymbol{\mu}_{a}$, $\boldsymbol{\mu}_{b}$ and the factors $\boldsymbol{L}_{11}$ and $\boldsymbol{L}_{21}$ of $\boldsymbol{\Lambda}_{aa}$ and $\boldsymbol{\Lambda}_{ab}$ coming from the Cholesky decomposition of $\boldsymbol{\Lambda}$. In short, the inverse covariance trick not only saves memory and computation but also reduces the number of optimizable parameters. Please see Sec. 4.3 for a benchmark comparison and Sec. A.1 for more details.
Figure 5: Benchmark Results on conditioning for 14’876 HyperGaussians with conditional attribute dimension $m=3$ (e.g., position) and varying latent dimension $n$. We average the measurements across 1000 runs after initial warm-up. The point measurements were conducted on an NVIDIA GeForce RTX 2080 Ti 11GB. The benchmark code performs one forward and one backward pass through the HyperGaussian module, which outputs the conditional mean and the uncertainty.
3.3 FlashAvatar with HyperGaussians
Our HyperGaussian representation can be integrated into existing 3DGS pipelines. To demonstrate this, we inject our conditional HyperGaussians into FlashAvatar [51]. FlashAvatar deploys a deformation MLP $F_{\theta}$, which maps FLAME [27] expression parameters $\boldsymbol{\psi}$ to per-Gaussian offsets $\Delta\boldsymbol{\mu}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{r}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{s}_{\boldsymbol{\psi}}$ for position, rotation, and scale. These offsets are then applied to the mesh-attached Gaussians to help model fine details. The deformation MLP further relies on auxiliary input consisting of positional encodings [34] of canonical mesh positions $\boldsymbol{\mu}_{T}$.
Fig. 2 illustrates how we replace 3DGS with HyperGaussians. We modify the deformation MLP to output a per-Gaussian latent $\boldsymbol{z}_{\boldsymbol{\psi}}$ instead of offsets. We then compute the conditional distributions $p(\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}})$, $p(\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}})$ and $p(\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}})$ for every Gaussian. Recall from Sec. 3.2 that the conditional means and covariance matrices can be computed efficiently in closed form. The conditional means are then fed to the rest of the pipeline analogous to the offsets $\Delta\boldsymbol{\mu}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{r}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{s}_{\boldsymbol{\psi}}$ from FlashAvatar. More specifically, the parameters in global space are obtained via
$\begin{split}\boldsymbol{\mu}_{\boldsymbol{\psi}}&=\boldsymbol{\mu}_{M}\oplus\mathbb{E}\bigl{[}\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}}\bigr{]}\\ \boldsymbol{r}_{\boldsymbol{\psi}}&=\boldsymbol{r}\oplus\mathbb{E}\bigl{[}\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}}\bigr{]}\\ \boldsymbol{s}_{\boldsymbol{\psi}}&=\boldsymbol{s}\oplus\mathbb{E}\bigl{[}\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}}\bigr{]},\end{split}$ (7)
where $\boldsymbol{\mu}_{\boldsymbol{\psi}}$ are the positions on the posed mesh and $\boldsymbol{r}$, $\boldsymbol{s}$ are base rotation and scale. We find a latent dimensionality of $n=8$ to perform best for face avatars, but already a single latent dimension ($n=1$) improves over the baseline in Tab. 3.
Our method improves the final quality for thin structures like the glass frames, teeth, and specular reflection in Fig. 1 and Fig. 7. For further details about FlashAvatar, please refer to Xiang et al. [51].
Figure 6: Uncertainty Quantification (Sec. 4.2) on one of the training subjects. Green denotes low uncertainty, while Red denotes high uncertainty. Note that the semantic structure arises purely from the formulation without additional supervision.
Figure 7: Qualitative Comparison with FlashAvatar [51], MonoGaussianAvatar [8], and SplattingAvatar [46]. Ours achieves high-quality details for thin structures (glass frames and teeth in the top row), specular reflections (eyes in the third row), and gracefully handles complex deformations (mouth in the second and fourth row).
Figure 8: Cross-reenactment Comparison with FlashAvatar [51], MonoGaussianAvatar [8], and SplattingAvatar [46]. Ours preserves fine details in the teeth and the overall shape of the subject. Please see the supplementary HTML page for more cross-reenactment results.
4 Experiments
This section outlines the experimental setting in Sec. 4.1, compares with the state-of-the-art for face avatars in Sec. 4.2, and ablates the key components of HyperGaussian in an ablation study in Sec. 4.3.
4.1 Setting
Dataset
We compare face avatars trained on subjects from $4$ datasets. For quantitative results in Tab. 2 and qualitative comparisons in Figs. 7 and 8, we compare on 5 datasets from previous works [61, 58, 14, 12, 11], consisting of a total of 19 subjects. All videos are sub-sampled to 25 FPS and resized to a resolution of $512\times 512$. The length of the videos varies between $1$ and $3$ minutes, and we train on $2000$ frames and use the last $500$ frames for testing, following related works [51].
Preprocessing
For preprocessing, we use the same pipeline as FlashAvatar, which consists of MICA [62] for FLAME tracking, RVM [28] for foreground matting, and an off-the-shelf face parser based on BiSeNet [55] for segmentation of the head and neck region, as well as the mouth.
Training Details
Our case study on FlashAvatar [51] builds directly on top of their public PyTorch [40] codebase. We inherit all losses and hyperparameters as shown in Sec. 3.3 and set the learning rate for our HyperGaussian parameters to $\eta=10^{-4}$. For each video, we train for $00030^{\prime}$ 000 iterations, which results in roughly $10$ epochs and takes between $15$ and $20$ minutes on a 24GB NVIDIA RTX 4090.
4.2 Comparisons
We compare with the state-of-the-art for Gaussian-based face avatars: MonoGaussianAvatar [8], SplattingAvatar [46], and FlashAvatar [51]. Tab. 2 shows the quantitative comparison against the state-of-the-art on the commonly used metrics PSNR, SSIM, and LPIPS [57] with a VGG [47] backbone.
Fig. 7 shows a qualitative comparison for self-reenactment. SplattingAvatar [46] is not designed to represent expression-dependent appearance effects. This leads to unrealistic renderings for reflective regions like eyes and glasses, and artifacts in regions with strong deformations like the mouth and the eyes. Note how it fails to close the eyelid in the second row. MonoGaussianAvatar [8] predicts expression-dependent offsets but struggles with specular reflections. FlashAvatar uses an MLP to predict expression-dependent Gaussian parameter offsets to optimized 3D Gaussians. Their output lacks detail for thin structures like the glass frames or gaps between teeth, and it does not work well for specular reflections on the eyes and glasses. Ours handles such high-frequency details more gracefully. Keep in mind that the only difference between Ours and FlashAvatar is the Gaussian representation. Without any other changes to the architecture and no hyperparameter tuning, HyperGaussian is capable of rendering more accurate specular reflections and thin structures in the mouth, eyes, and glass frames.
We qualitatively compare cross-reenactment in Fig. 8. Note the distorted mouth and blurry hair in SplattingAvatar. MonoGaussianAvatar renders a more realistic face but exhibits unrealistic deformations on the jaw. FlashAvatar can render the correct expression, but the teeth and hair appear blurry. Ours produces higher-quality renders showing individual teeth and sharper hair. This can also be observed in a side-by-side comparison for training convergence in Fig. 3. Our HyperGaussian leads to a faster conversion than vanilla 3DGS.
Finally, it’s important to mention that MonoGaussianAvatar is substantially heavier than FlashAvatar and Ours. MonoGaussianAvatar requires over 12 hours of training time and uses over 100,000 Gaussians. FlashAvatar and Ours are much more lightweight with only $14,876$ Gaussians and less than 20 minutes training time. Despite the much longer training and the higher number of Gaussians for related works, our HyperGaussians outperforms them both. We encourage the reader to check out the supplementary HTML page to see animated results.
Method PSNR ↑ SSIM$\bigl{(}10^{-1}\bigr{)}$ LPIPS$\bigl{(}10^{-2}\bigr{)}$
SplattingAvatar [46] 28.58 9.396 9.021
MonoGaussianAvatar [8] 29.94 9.456 6.545
FlashAvatar [51] 29.43 9.466 5.107
Ours 29.99 9.510 4.978
Table 2: Quantitative comparison with state-of-the-art digital avatar reconstruction methods from monocular video across 19 subjects from 5 datasets [61, 58, 14, 12, 11]. Ours corresponds to FlashAvatar with 8-dimensional HyperGaussians without any other modifications. Please see Sec. 3.3 for details.
Uncertainty from Covariances
We observe an emerging property about HyperGaussians, which arises naturally throughout training. HyperGaussians are at their core multivariate Gaussian distributions. Their conditional covariance matrices indicate the variance of each Gaussian across the different expressions of the training subject and can be intuitively interpreted as uncertainty. We compute the uncertainties $\sigma:=\log\det\boldsymbol{\Sigma}_{a|b}=-2\operatorname{tr}\log\boldsymbol{L}_{aa}$ for each conditional distribution $p(\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}})$, $p(\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}})$, and $p(\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}})$. Please refer to Eq. 14 in the Sec. A.1 for details. We visualize an example in Fig. 6. Red indicates high uncertainty, and green indicates low uncertainty. This agreement between the uncertainty estimates and what would intuitively be considered difficult regions is an inductive bias of our formulation and does not require explicit supervision.
4.3 Ablation Study
Inverse Covariance Trick
As discussed in Sec. 3.2, a naive implementation of the conditioning in Eq. 4 is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$ because it requires inverting the large matrix $\boldsymbol{\Sigma}_{bb}$. The bottleneck lies in storing the covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$. After applying the inverse covariance trick (Sec. 3.2), we only need to construct $\boldsymbol{\Sigma}_{aa}\in\mathbb{R}^{n\times n}$ by inverting $\boldsymbol{\Lambda}_{aa}^{-1}$, a small matrix. As an example, the matrix for the position attribute is only $\boldsymbol{\Lambda}_{aa}^{-1}\in\mathbb{R}^{3\times 3}$. Fig. 5 shows an empirical ablation for latent dimensionalities $n$ between $1$ and $128$. In our case study on FlashAvatar [51] (Sec. 3.3), the inverse covariance trick improves speed by $150\%$ for a small latent with $n=8$ and by $15,000\%$ for a large latent with $n=128$. Not only does the inverse covariance matrix improve speed, but it also reduces memory usage. For a small latent with $n=8$, the naïve implementation uses $42$ MB, whereas the inverse covariance trick reduces this to $22$ MB, a reduction of $48\%$. For the large latent with $n=128$, the reduction is over 90%.
Latent Dimensionality
We ablate the effect of different latent dimensionalities ($n$ in Sec. 3.2) in Tab. 3. We find that HyperGaussian are robust towards different latent dimensions. A latent dimension of $8$ performs very well, but we already observe an improvement for a single latent dimension ($n=1$) over the vanilla 3DGS variant. The bottom row corresponds to the FlashAvatar baseline [51], which doesn’t use any HyperGaussians.
Latent dim. ($n$) PSNR ↑ SSIM$\bigl{(}10^{-1}\bigr{)}$ LPIPS$\bigl{(}10^{-2}\bigr{)}$
1 29.73 9.492 5.066
2 29.92 9.503 5.000
4 29.89 9.507 4.994
8 29.99 9.510 4.978
16 29.92 9.511 4.978
32 29.89 9.511 4.976
64 29.91 9.512 5.020
Baseline 29.43 9.466 5.107
Table 3: Quantitative Ablations. Green denotes the best and Yellow the second best. The bottom row indicates the FlashAvatar [51] baseline, which does not use any HyperGaussians. We train the same model with different latent dimensions, ranging from $n=1$ to $n=64$. Employing HyperGaussians with a single latent dimension ($n=1$) already outperforms the baseline. The differences above $8$ dimensions ($n=8$) are minor, which indicates that 8 dimensions are sufficiently expressive.
5 Conclusion and Discussion
In this paper, we study how 3D Gaussian Splatting can be made more expressive for monocular face avatars. The result, HyperGaussians, is a novel extension to 3D Gaussians that offers improved expressivity and rendering quality, excelling in rendering high-frequency details like thin structures and specular reflections. Our evaluations on 19 subjects from 4 datasets outperform the state-of-the-art by simply plugging in our proposed HyperGaussians into an existing method, FlashAvatar, without any other modifications. As a limitation, HyperGaussians are high-dimensional representations and require more memory than vanilla 3DGS. Our proposed inverse covariance trick greatly reduces these requirements, but scaling HyperGaussian to thousands of dimensions is a topic for future work. In addition, HyperGaussians are not a stand-alone method, and hence, they inherit limitations of the underlying method. Future research directions could investigate the effect of HyperGaussians beyond human faces, e.g., full-body avatars or more generic dynamic scenes. In conclusion, HyperGaussians show great promise in improving high-frequency details for monocular face avatars, bringing the field a step closer to photorealistic and fast monocular face avatars.
Acknowledgments
We thank Jenny Schmalfuss, Seonwook Park, Lixin Xue, Zetong Zhang, Chengwei Zheng, Egor Zakharov for fruitful discussions and proofreading, and Yufeng Zheng for help with dataset preprocessing.
Appendix A Supplementary
This supplement contains more details and derivations in Sec. A.1, supplementary results in Sec. A.2, and discusses the potential societal impact of this work in Sec. A.3.
A.1 HyperGaussian Details
A.1.1 Formulation
This section provides more details about the formulation of HyperGaussian. We first explain how we parameterize the mean and covariance matrix. We then detail how we construct the covariance for splatting and derive the inverse covariance trick in more detail. Finally, we explain how we extract an uncertainty measure from the optimized covariance matrices.
Parameterization
Each HyperGaussian consists of a (typically high) dimensional mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$ with optimizable parameters. We parameterize the mean $\boldsymbol{\mu}$ directly and decompose the covariance $\boldsymbol{\Sigma}$ into its Cholesky factor [10] $\boldsymbol{L}$, such that $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}$, where $\boldsymbol{L}$ is a lower triangular matrix with positive diagonal entries. To ensure the uniqueness of the factorization, we apply an exponential activation function, $\boldsymbol{L}_{i,i}(x)=e^{x}$, to the diagonal entries of the parameter matrix.
Splatting Details
This paragraph explains how we construct the covariance matrix for splatting the conditioned Gaussians [7]. It is important to note that the HyperGaussian conditional covariance $\boldsymbol{\Sigma}_{a|b}$ is not equal to the covariance $\boldsymbol{\Sigma}$ from Eq. 1. The conditional covariance $\boldsymbol{\Sigma}_{a|b}$ is used to compute the conditional mean $\boldsymbol{\mu}_{a|b}$. The covariance used for splatting is constructed from the conditional means for rotation and scale:
$\boldsymbol{\mu}_{R|\gamma_{b}}\boldsymbol{\mu}_{S|\gamma_{b}}\boldsymbol{\mu}_{S|\gamma_{b}}^{\top}\boldsymbol{\mu}_{R|\gamma_{b}}^{\top}.$ (8)
The conditional covariance $\boldsymbol{\Sigma}_{a|b}$ is not actually used for splatting. We only use it for visualizing the uncertainties (see Sec. A.1.1). This is an important distinction to NDGS [10], which directly splats using the conditional matrix. NDGS has no degrees of freedom for the conditional orientation of the 3D Gaussians, i.e., the 3D Gaussians cannot rotate.
Derivation of the Inverse Covariance Trick
We explain in the main paper that a naïve implementation of the conditioning is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$. Here, we provide a more detailed derivation of the covariance trick.
The bottleneck lies in storing and inverting the conditional covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$, which tends to be a large matrix. Ideally, we would like to transfer the inversion of $\boldsymbol{\Sigma}_{bb}$ to the smaller and constant-sized block matrix $\boldsymbol{\Sigma}_{aa}$. To achieve this, we start by reformulating our HyperGaussians in terms of their precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ such that $\boldsymbol{\gamma}\sim\mathcal{N}\bigl{(}\boldsymbol{\mu},\boldsymbol{\Lambda}^{-1}\bigr{)}$. We again consider the general block matrix view:
$\boldsymbol{\Sigma}^{-1}=\boldsymbol{\Lambda}=\begin{bmatrix}\boldsymbol{\Lambda}_{aa}&\boldsymbol{\Lambda}_{ab}\\ \boldsymbol{\Lambda}_{ba}&\boldsymbol{\Lambda}_{bb}\end{bmatrix}$ (9)
with $\boldsymbol{\Lambda}_{ba}=\boldsymbol{\Lambda}_{ab}^{\top}$.
As the inverse of $\boldsymbol{\Sigma}$, $\boldsymbol{\Lambda}$ inherits symmetry, Eq. 10, as well as positive definiteness since its eigenvalues are the reciprocal of the eigenvalues of $\boldsymbol{\Sigma}$, and therefore all positive:
$\boldsymbol{\Lambda}^{\top}=\bigl{(}\boldsymbol{\Sigma}^{-1}\bigr{)}^{\top}=\bigl{(}\boldsymbol{\Sigma}^{\top}\bigr{)}^{-1}=\boldsymbol{\Sigma}^{-1}=\boldsymbol{\Lambda}.$ (10)
This is important, as it allows us to reuse the same parameterization that we described in Sec. A.1.1 to represent $\boldsymbol{\Lambda}=\boldsymbol{L}\boldsymbol{L}^{\top}$. Conveniently, we also get the Cholesky factor $\boldsymbol{L}_{11}$ of $\boldsymbol{\Lambda}_{aa}$ as a side product:
$\begin{split}\begin{bmatrix}\boldsymbol{\Lambda}_{aa}&\boldsymbol{\Lambda}_{ab}\\ \boldsymbol{\Lambda}_{ba}&\boldsymbol{\Lambda}_{bb}\end{bmatrix}&=\begin{bmatrix}\boldsymbol{L}_{11}&\boldsymbol{0}\\ \boldsymbol{L}_{21}&\boldsymbol{L}_{22}\end{bmatrix}\begin{bmatrix}\boldsymbol{L}_{11}^{\top}&\boldsymbol{L}_{21}^{\top}\\ \boldsymbol{0}&\boldsymbol{L}_{22}^{\top}\end{bmatrix}\\ &=\begin{bmatrix}\boldsymbol{L}_{11}\boldsymbol{L}_{11}^{\top}&\boldsymbol{L}_{11}\boldsymbol{L}_{21}^{\top}\\ \boldsymbol{L}_{21}\boldsymbol{L}_{11}^{\top}&\boldsymbol{L}_{21}\boldsymbol{L}_{21}^{\top}+\boldsymbol{L}_{22}\boldsymbol{L}_{22}^{\top}\end{bmatrix}.\end{split}$ (11)
With this new formulation, the conditional mean and covariance matrix can be expressed as
$\begin{split}\boldsymbol{\mu}_{a|b}&=\boldsymbol{\mu}_{a}-\boldsymbol{\Lambda}_{aa}^{-1}\boldsymbol{\Lambda}_{ab}\bigl{(}\boldsymbol{\gamma}_{b}-\boldsymbol{\mu}_{b}\bigr{)}\\ \boldsymbol{\Sigma}_{a|b}&=\boldsymbol{\Lambda}_{aa}^{-1},\end{split}$ (12)
where the term $\boldsymbol{\Lambda}_{aa}^{-1}\boldsymbol{\Lambda}_{ab}$ can be further broken down into
$\boldsymbol{\Lambda}_{aa}^{-1}\boldsymbol{\Lambda}_{ab}=\bigl{(}\boldsymbol{L}_{11}\boldsymbol{L}_{11}^{\top}\bigr{)}^{-1}\boldsymbol{L}_{11}\boldsymbol{L}_{21}^{\top}=\boldsymbol{L}_{11}^{-\top}\boldsymbol{L}_{21}^{\top}.$ (13)
Note that products with $\boldsymbol{L}_{11}^{-\top}$ can be evaluated in an efficient and numerically stable manner since $\boldsymbol{L}_{11}$ is a triangular matrix of small size, which is independent of the latent dimension. Please see the main paper for a benchmark comparison between the naïve implementation and the one applying the inverse covariance trick.
Derivation of Uncertainty
As discussed in the main paper, we observe an interesting property about HyperGaussians, which arises naturally throughout training. HyperGaussians are at their core multivariate Gaussian distributions. Their conditional covariance matrices indicate the variance of each Gaussian across the different expressions of the training subject and can be intuitively interpreted as uncertainty.
More formally defined, we have
[…]
where we used $\det\boldsymbol{\Sigma}_{a|b}=\det\boldsymbol{\Lambda}_{aa}^{-1}=\bigl{(}\det\boldsymbol{\Lambda}_{aa}\bigr{)}^{-1}$ in step $(1)$, $\det\boldsymbol{\Lambda}_{aa}=\det\boldsymbol{L}_{aa}\boldsymbol{L}_{aa}^{\top}=\bigl{(}\det\boldsymbol{L}_{aa}\bigr{)}^{2}$ in step $(2)$, and $\det\boldsymbol{L}_{aa}=\prod_{i=1}^{m}(\boldsymbol{L}_{aa})_{i,i}$ in step $(3)$. The $\log$ at step $(3)$ is applied element-wise.
Again, we can compute this quantity efficiently using the inverse covariance trick (Sec. A.1.1). These values are summed up across the conditional distributions for position, rotation, and scale. In order to render these uncertainties, we further apply a sigmoid function and map the values to colors. This agreement between the uncertainty estimates and what would intuitively be considered difficult regions emerges without explicit supervision.
A.2 Supplementary Experiments
A.2.1 Qualitative Results
We show supplementary video results for self- and cross-reenactment on a supplementary HTML page.
FlashAvatar
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FlashAvatar
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
5 seconds 20 seconds 60 seconds 180 seconds 600 seconds
Figure 9: We compare the convergence speed of FlashAvatar [51] vs. Ours. The only difference between FlashAvatar and Ours is the substitution of 3D Gaussians (top) with HyperGaussians (bottom), as described in the case study in the main paper. From the beginning, HyperGaussians display sharper results.
In addition, Fig. 9 provides more examples for comparing the convergence speed of FlashAvatar without (top) vs. with HyperGaussians (bottom).
A.2.2 Ablation Study
We complement the ablation study from the main paper with supplementary results for different MLP configurations in Tab. 4 and Fig. 12. The default FlashAvatar MLP [51] has 6 layers with 256 neurons, totaling 375K parameters. Replacing vanilla 3D Gaussians with HyperGaussians adds optimizable parameters (see Sec. A.1.1). One might assume that simply increasing the parameter count for the FlashAvatar MLP would improve the results, but this is not the case. We ablate different MLP configurations in Tab. 4. Adding more parameters to the MLP does not perform as well as adding HyperGaussian. In fact, it performs the same while slowing down the rendering speed. FlashAvatar with vanilla 3DGS runs at 347 FPS. With a large MLP, this number drops to 158 (for $256\times 40$, 2.6 M parameters) and 178 (for $512\times 11$, 2.7 M parameters). With HyperGaussian ($n=8$ and 2.6 M parameters), the original MLP ($256\times 6$) outperforms the other MLP variants for all metrics while maintaining a rendering speed of 300 FPS. All metrics and rendering times were computed on a single Nvidia GeForce RTX 2080 Ti for images with resolution $512\times 512$. In summary, the HyperGaussians’ performance improvement cannot be matched by increasing the complexity of the MLP. HyperGaussians boost the performance while maintaining fast rendering speed.
A fundamental advantage of HyperGaussians is their ability to distill highly local context, enabling independent deformations between spatially proximate but semantically distinct regions. For instance, our method can independently model glass frames near the upper cheek or the upper teeth adjacent to the jaw. In contrast, FlashAvatar suffers from stronger coupling between neighboring Gaussians due to its shared MLP architecture and direct offset approach. This coupling creates an optimization challenge where improvements in one region often degrade quality in others. Our approach allows each region to optimize independently, preserving detailed geometry and appearance across semantically different but spatially adjacent facial features.
Figure 10: Comparison with GaussianHeadAvatar on multiview videos from the NeRSemble [19] dataset. HyperGaussians (Ours) demonstrate more accurate reconstructions for complex deformations, thin structures, and specular highlights. We outperform in all metrics: higher PSNR, SSIM ($10^{-1}$), and lower LPIPS ($10^{-1}$).
A.2.3 GaussianHeadAvatar Integration
We demonstrate the broad applicability of HyperGaussians by successfully integrating them into GaussianHeadAvatar [53]. Fig. 10 shows significant visual improvements: more accurate modeling of complex skin deformations (rows 1, 2), better alignment of fine structures in hair strands (row 2), and sharper specular reflections in the eyes (row 1). These quality enhancements come with minimal computational overhead, increasing training time by only 5.6% (from 9h to 9.5h). These results confirm that HyperGaussians can effectively enhance existing Gaussian-based avatars with negligible performance impact.
Figure 11: Qualitative comparison with NDGS integrated into FlashAvatar. The limited degrees of freedom of NDGS cause misalignments of thin structures and edges. The numbers show PSNR, SSIM ($10^{-1}$), and LPIPS ($10^{-2}$).
A.2.4 Comparison with NDGS
Our approach fundamentally differs from NDGS [10] in its mathematical formulation and capabilities. While NDGS uses the conditional covariance matrix, which is independent of the latent code, to directly represent the size and shape, our HyperGaussians apply multivariate Gaussians on each attribute where the conditional means dynamically adapt the location, scale, and orientation of the derived 3D Gaussians in response to the latent code. This crucial difference gives HyperGaussians the necessary degrees of freedom to model thin geometry and complex deformations with higher accuracy. Fig. 11 shows our method eliminates the artifacts visible in NDGS (particularly on glass frames) and produces substantially more precise geometry at boundaries.
Figure 12: Qualitative comparison for varying latent dimensionalities. We find that HyperGaussians are robust towards different latent dimensions. A latent dimension of $8$ performs best, but we already observe an improvement for a single latent dimension ($n=1$) over the vanilla 3DGS variant.
MLP (width×$\times$×depth) HyperGaussian? Latent # Param. PSNR ↑ SSIM $\bigl{(}10^{-1}\bigr{)}$ LPIPS$\bigl{(}10^{-2}\bigr{)}$
256$\times$6 no - 375 K 29.43 9.466 5.107
256$\times$40 no - 2.6 M 28.10 9.380 5.720
512$\times$11 no - 2.7 M 29.50 9.472 5.122
256$\times$6 yes 1D 1.2 M 29.73 9.492 5.066
256$\times$6 yes 2D 1.4 M 29.92 9.503 5.000
256$\times$6 yes 4D 1.8 M 29.89 9.507 4.994
256$\times$6 yes 8D 2.6 M 29.99 9.510 4.978
256$\times$6 yes 16D 4.1 M 29.92 9.511 4.978
256$\times$6 yes 32D 7.2 M 29.89 9.511 4.976
256$\times$6 yes 64D 13.4 M 29.91 9.512 5.020
Table 4: Supplementary Ablations. Simply increasing the parameter count for the FlashAvatar MLP does not improve the metrics. Our HyperGaussians, however, improve the performance of the original MLP out-of-the-box. As an additional benefit, HyperGaussians render 300 FPS while the deeper MLPs drop to 158 FPS ($256\times 40$) and 178 FPS ($512\times 11$), respectively. FlashAvatar with vanilla 3DGS renders fastest at 347 FPS. Green denotes the best and Yellow the second best.
A.3 Societal Impact
It is important to be aware that photorealistic, high-quality face avatars from monocular videos can have societal implications. While our novel HyperGaussian representation contributes to exciting possibilities for entertainment, communication, and virtual experiences, it could potentially be misused to spread misinformation and deception. Realistic face avatars could be exploited to produce convincing deepfakes, potentially undermining trust in visual media and influencing societies and politics. We strongly condemn any form of abuse or malicious use of our research and advocate for responsible development and application of face avatar technology, always in strict accordance with local laws and regulations.